# Fixed-point Inversion for Text-to-image diffusion models

### v1

@@ 3.3 Fixed-point iterations for consistent inversion
This problem is illustrated in Figure 3(a). Here, we randomly sampled $z_{T}$ and ran the forward process with the prompt “A cute dog”. This results in a denoised latent $z_{0}$ which is decoded into an image $x$. Then, $z^{E}_{0}$ is obtained by passing $x$ through the encoder $E$. In terms of $L_{2}$ distance, $z^{E}_{0}$ differs only slightly from $z_{0}$, but it is not consistent with the prompt $p$ (outside the dashed circle). As a result, applying inversion and reconstruction (dashed brown lines) from $z^{E}_{0}$ yields a reconstructed latent $\hat{z}^{E}_{0}$ that is far from $z_{0}$. As a consequence, when $z^{E}_{0}$ is decoded by the VAE, it results in a different image $x^{\prime}$ of “a cute dog”. This illustrates that inversion becomes inconsistent when applied to prompt-agnostic latents. To quantify this effect, we repeated this experiment for 10k images and found the mean $L_{2}$ distance between $z_{0}$ and $\hat{z}^{E}_{0}$ ($||z_{0}-\hat{z}^{E}_{0}||$) to be $99$. In comparison, the mean distance between $z_{0}$ and $z^{E}_{0}$ ($||z_{0}-z^{E}_{0}||$) is only $16$. This indicates that the inversion process produces latents that deviate significantly from the original $z_{0}$.
How can this inconsistency be resolved? We propose to employ a short preprocessing step of another fixed-point iteration. The objective is to find a fixed point $z_{0}$ for the implicit function $z_{0}=h(z_{0},p)$, with $h$ representing the backward-forward process and $p$ is the prompt that would be used for inversion. The initial guess for the fixed-point iterations is $z^{E}_{0}=E(x)$. The iteration count should be kept small to prevent excessive alteration of the latent $z_{0}$ as this could potentially affect the visual appearance of the original image. This process is also known as Image2Image editing [11]. In practical terms, we observed that a brief cycle of fixed-point iteration, involving two noise-addition steps (backward process) followed by two denoising steps with the prompt $p$ (forward process), proves effective.
This process is illustrated in Figure 3(b). Here, we pre-process $z^{E}_{0}$ by two fixed-point iterations to obtain an improved latent ${\tilde{z}}_{0}$ (red lines). ${\tilde{z}}_{0}$ is decoded to a very similar image $x$ and results with consistent inversion (dashed green lines). For 10k images, we found that the $L_{2}$ distance between $z_{0}$ and ${\tilde{z}}_{0}$ (i.e., $||z_{0}-{\tilde{z}}_{0}||$) is only $12$. This illustrates that the proposed pre-processing method is capable of maintaining a highly consistent backward-forward process.
@@
Figure 3: Prompt-aware adjustment for more consistent inversion. (a) A seed $z_{T}$ is used to generate latent image $z_{0}$, which is then decoded into an image $x$. If that image is encoded into the latent space $z^{E}_{0}=E(x)$, its representation $z^{E}_{0}$ may not be aligned with the prompt used for inversion and generation (brown dashed line). As a result, the reconstructed $\hat{z}^{E}_{0}$ differs significantly from $z_{0}^{E}$, yielding an inconsistent image $x^{\prime}$. (b): Inconsistency can be fixed using a short noising and denoising step (red curves), which yields $\tilde{z}_{0}$. Then, applying the complete backward-forward process (green curves) can generate a consistent $\tilde{z}_{0}$.
@@ 4 Experiments
We evaluate our approach on three main tasks: (1) Image inversion and reconstruction: Here, We assess the inversion fidelity by evaluating the quality of the reconstructed images. (2) Image Editing: We show image editing performance using our inversion scheme and, (3) Rare Concept Generation: We illustrate how our method can be applied to improve the generation of rare concepts. Code will be made publicly available upon acceptance.
Compared Methods: We compared our approach with the standard DDIM Inversion [21] and with two state-of-the-art inversion techniques: (1) Null-text [13] which optimizes the null conditioning at every denoising step to accurately invert a given image and, (2) EDICT [22], an inversion method based on affine coupling layers. In all experiments, we used code and parameters provided by the authors of the compared methods.
Implementation details:
All experiments were conducted with Stable Diffusion v1.5 [16] implemented with Pytorch. Input images were trimmed to make them square and resized to 512 × 512. For all approaches, we set the number of sampling steps to be 50 with a stride of 20 over $T=1000$ diffusion steps. All baselines were run until convergence, while FPI was run for fixed 5 iterations per diffusion step. For a fair comparison, all methods were executed on the same infrastructure (a single A100 GPU), and the experiment runtime was documented accordingly.
Figure 4: Inversion Results: Mean reconstruction quality (y-axis, PSNR) and runtime (x-axis, seconds) of four inversion methods on the COCO2017 validation set. All methods run for 50 denoising steps. Our method reaches high PSNR with ×$\times 4$ to ×$\times 12$ shorter inversion-reconstruction time compared to other methods. The dashed line represents an upper bound induced by the VAE. All runtimes are measured on a single A100 GPU.
@@ 4.1 Image Reconstruction
To evaluate the fidelity of our approach, we measure PSNR of images reconstructed from seed inversions. Specifically, we used the entire set of 5000 images from the MS-COCO-2017 validation dataset [9], along with their corresponding captions. For each image-caption pair, we first found the inverted latent and then used it to reconstruct the image using the same caption. Given that the COCO dataset provides multiple captions for each image, we used the first listed caption as a conditioning prompt.
Figure 4 shows PSNR of reconstructed images in relation to inversion time, demonstrating the performance of our approach compared to state-of-the-art (SoTA) inversion methods. Time is measured on a single NVIDIA A100 GPU for all methods. The dashed black line is the upper bound induced by the Stable Diffusion VAE. It shows that our method is able to achieve comparable PSNR to recent methods and close to the upper bound of VAE, yet accomplishes this in the shortest amount of time. Furthermore, in contrast to other methods, our approach accurately inverts an image to a single latent vector without requiring additional memory. Figures LABEL:fig:editing and 5 further demonstrate a qualitative comparison of reconstructed images. We provide in the supplementary material an additional analysis.
Memory Usage: Compared to Null-text [13] and EDICT [22], FPI does not require any additional learnable parameters. In contrast, EDICT requires 16K additional parameters, and and Null text necessitates almost 3 million additional parameters for a single image.
Figure 5: Image reconstruction, qualitative results : We show reconstructions of images from MS-COCO, focusing on areas with fine details. At the top row, FPI correctly reconstructs the eye and fur of the dog and at the bottom row, it correctly reconstructs the label text on the bottle (Mikkeller). All baselines were executed until they achieved convergence.
Figure 6: Qualitative results of image editing. FPI edits images more naturally while preserving the structure of the original image. All baselines were executed until they reached convergence, whereas our approach was run for five iterations per diffusion step.
4.2 Image Editing
@@
We now evaluate the effect of inverting images with FPI, compared to state-of-the-art inversion baselines, for complex real image editing. We used Prompt2Prompt [6] as our editing method. All baselines were executed until they reached convergence, whereas our approach was iterated for a fixed number of five iterations per diffusion step (editing in 5 seconds).
We present below both qualitative and quantitative results.
@@
Qualitative results. Figure LABEL:fig:editing and 6 give a qualitative comparison between Null-text, EDICT, and FPI in the context of real image editing. FPI excels in accurately editing real images with target prompts producing images with both high fidelity to the input image and adherence to the target prompt. The examples illustrate how alternative approaches may struggle to retain the original structure or tend to overlook crucial components specified in the target prompt. For instance, in Figure LABEL:fig:editing(right), competing methods largely fail to add the boat as requested. In row 2 of Figure 6, FPI transforms only the bananas into oranges, preserving the appearance of the woman, that hat, and people in the background. In the third row, Null-text [13] eliminates the woman behind the man, and EDICT [22] removes the pile of apples on the left. FPI stands out in altering only the fruits without affecting the rest of the image. More examples can be found in the supplementary material.
Null Text EDICT FPI (ours)
Percent preferred 18.3% 20.0% 61.7%
Table 1: User Study: Values are the percentage of raters that selected each option. N=60 images, of which 12 were provided by the authors of [13] and the rest were randomly sampled from COCO. Raters were asked to select the best editing result in terms of fidelity to both the input image and to an edit instruction.
Figure 7: Semantic quality vs reconstruction quality: Four inversion methods are compared. Semantic quality is measured using CLIP distance between image and text prompt. The image reconstruction quality is measured using LPIPS distortion score following previous protocols.
ImageNet1k in LAION2B
Many Med Few Total Acc FID $\hat{T}_{Init}$ $\hat{T}_{Opt}$
n=235 n=509 n=256 (sec) (sec)
Methods #>1M 1M>#>10K 10K>#
NAO w DDIM 98.5 96.9 85.1 94.3 6.4 25 29
NAO w FPI (ours) 98.6 97.2 87.3 95.0 6.4 19 26
Table 2: Impact of inversion quality on the generation of rare concepts: Image generation quality measured by the accuracy of a pre-trained classifier. We compare NAO [18], a state-of-the-art algorithm when using DDIM inversion (top row), and our FP-inversion (bottom row). Average per-class accuracy is reported separately for the Head (over 1M samples in train set), Medium, and Tail (rare classes below 10K train samples). Our FPI boosts the accuracy of rare and medium concept generation without jeopardizing performance elsewhere.
@@
User Study. We further evaluated editing quality using human raters. We followed the evaluation process of [13] and asked human raters to rank edits made by three methods. 60 images were rated, 12 images were provided by the authors of [13] and the rest were randomly selected from the COCO dataset [9]. Three raters for each image were recruited through Amazon Mechanical Turk and were tasked to choose the image that better applies the requested text edit while preserving most of the original image. As shown in Table 1, FPI was favored by raters by a large margin.
@@
FPI (ours) 54.98 5.91 70.18 4.59
Table 3: Impact of inversion on image interpolation and centroid finding. (NOA-Path and NOA-centroid [18]). In the interpolation task, two images $x^{1},x^{2}$ are inverted, and then images are generated along the path between their latent seeds $z_{T}^{1},z_{T}^{2}$. In centroid-finding, a set of images is inverted and their centroid is found. We measure the quality of generated images by NAO [18] where the only difference is the inversion methods. Acc and FID scores improved considerably when replacing DDIM with FPI.
LPIPS vs CLIP Score. Following [22, 8], we evaluate the results using two complementary metrics: LPIPS [24] to quantify the extent of structure preservation (lower is better) and a CLIP-based score to quantify how well the generated images comply with the text prompt (higher is better). Metrics are averaged across 100 MS-COCO images, (see Suppl. for details). Figure 7 illustrates that editing with FPI yields a superior CLIP and LPIPS score, demonstrating the ability to perform state-of-the-art real image editing with superior fidelity. It further affirms the findings derived from the user study.
4.3 Seed Interpolation & Rare concept generation
@@
Interpolation and centroid finding: We start by evaluating the benefit of our seed inversions for image interpolation and centroid finding. We follow the experimental protocol of [18] and compare images generated by DDIM with those produced by FPI. Evaluation is conducted based on FID score and image accuracy, assessed using a pre-trained classifier. See details in [18]. Results are presented in Table 3. Notably, in comparison to DDIM inversion, initializing with our seeds consistently results in higher-quality images both in interpolation paths and seed centroids. Visual illustration can be found in suppl. material.
@@
Rare concept generation: We further show the effect of our seed inversions on the performance of NAO centroids with SeedSelect for rare concept generation. Specifically, we compared images generated by SeedSelect initialized with NAO using DDIM inversion (NAO(DDIM) in Table 2) and using FPI (NAO(FPI)). We followed the evaluation protocol of [19, 18] on ImageNet1k classes arranged by their prevalence in the LAION2B dataset [20]. This dataset serves as a substantial “in the wild” dataset employed in training foundation diffusion models, such as Stable Diffusion [16]. Image quality is measured by FID and accuracy of a pre-trained classifier. For more details see [19, 18].
@@
The results, summarized in Table 2, demonstrate that our inversion method significantly boosts performance, both in accuracy and FID, compared to DDIM inversions. Furthermore, our inversions yield a more precise and effective initialization point for SeedSelect [19], resulting in notably quicker convergence (see Table 2) without compromising accuracy or image quality.
5 Summary
Image inversion in diffusion models is a crucial building block for various applications, including image editing, semantic augmentation, and generating images of rare concepts. The quality of the inverted seed in the noise space significantly affects downstream task performance. Current methods trade off inversion quality with computational cost (runtime, and memory footprint), taking an order of magnitude more compute for high-quality inversion. This paper introduces Fixed-Point Inversion (FPI), a novel method for diffusion inversion of real images that establishes a new point in this trade-off map. FPI revolves around the implicit equation governing the diffusion process, replacing the approximation step with an iterative process involving fixed-point iterations. This approach converges rapidly and demonstrates superior performance in terms of accuracy, execution time, and memory footprint.
We evaluate FPI extensively in three downstream tasks including a) Image inversion and reconstruction, b) Real image editing, and c) Rare concept generation. In all cases, we find that our solution produces improved results compared to the state of the art. This was further supported by a user study, indicating a preference for images edited by our method. Our solution is easy to implement and does not involve additional training, eliminating the need for computationally expensive tuning of the diffusion model.
Supplementary Material
Figure 8: Qualitative and Speed comparison for image inversion: The results showcase the speed of FPI compared to SoTA. While FPI completes reconstruction after 5 sec, other alternatives are deeply at their intermediate stages.
Figure 9: Qualitative and Speed comparison for image editing: FPI achieves faster and often more plausible editing results while retaining high fidelity to the original image. Column 1: The change to a ”Cat” is more realistic. Column 2: The giraffe is well structured. Column 3: While other methods struggle in this case, FPI conducts the editing with a subtle change by creating table ”legs” to the suitcase (note the gap between the legs) and making the suitcase top flat, thus changing it to a table without a noticeable change to the rest of the image (e.g. the flowers). Column 4: A case where all methods succeed. Please zoom in for better observation.
6 Additional Qualitative Results
In this section, we present additional qualitative comparisons involving FPI, Null-Text [13], and EDICT [22]. Figure Fixed-point Inversion for Text-to-image diffusion models provides insights into inversion-reconstruction, showcasing that FPI converges significantly faster than previous techniques when it comes to image inversion.
Figure 10: Qualitative comparison for image editing A: One can notice the ”under” and ”over” editing effects in many cases. Row 1: Null-text changes the food only at the top plate, EDICT looks to miss the change while FPI succeeds. Row 2: Null-text adds a new object to the scene, EDICT, removes the person at the back, and adds a red tint to the woman’s hair color. Rows 3 & 4: FPI creates more reliable edits. Row 5: We show a case where all methods struggle to make a plausible editing. Please zoom in for better observation.
Figure 11: Qualitative comparison for image editing B: Note the 3rd row (clean water), where only our FPI was capable of making the correct change.
Figure 12: Various editing with same input: Note the FPI capability in both subtle and extensive changes as one would expect from the particular prompt change.
Figure 13: Effect of prompt-aware adjustment on image inversion. Here we show the strong impact of our prompt-aware adjustment on the reconstruction results. Without the adjustment, the results are often unrealistic or far from the original image.

### v2

@@ 4.2 Prompt-aware adjustment for consistent inversion
We now focus on an important subtlety that greatly impacts inversion quality. Our inversion process is designed to agree with a generation process for a prompt $p$. However, inversion is typically applied to photos and images that are not generated by the model. Such an image $x$ is mapped into the latent $z_{0}$ in a way that is unaware of the prompt because the VAE encoder is prompt-agnostic. This means that the typical input to our inversion process may be inconsistent with the prompt we use. This problem is illustrated in Figure 4(a). Here, we randomly sampled $z_{T}$ and ran the forward process with the prompt “A cute dog". This results in a denoised latent $z_{0}$ which is decoded into an image $x$. Then, $z^{E}_{0}$ is obtained by passing $x$ through the encoder $E$. In terms of $L_{2}$ distance, $z^{E}_{0}$ differs only slightly from $z_{0}$, but it is not consistent with the prompt $p$ (outside the dashed circle). As a result, applying inversion and reconstruction (dashed brown lines) from $z^{E}_{0}$ yields a reconstructed latent $\hat{z}^{E}_{0}$ that is far from $z_{0}$. As a consequence, when $\hat{z}^{E}_{0}$ is decoded by the VAE, it results in a different image $x^{\prime}$ of “a cute dog". This illustrates that inversion becomes inconsistent when applied to prompt-agnostic latents. To quantify this effect, we repeated this experiment for 10k images and found the mean $L_{2}$ distance between $z_{0}$ and $\hat{z}^{E}_{0}$ ($||z_{0}-\hat{z}^{E}_{0}||$) to be $99$. In comparison, the mean distance between $z_{0}$ and $z^{E}_{0}$ ($||z_{0}-z^{E}_{0}||$) is only $16$. This indicates that the inversion process produces latents that deviate significantly from the original $z_{0}$.
@@
Figure 4: Prompt-aware adjustment for more consistent inversion. (a) A seed $z_{T}$ is used for generating a latent image $z_{0}$, then decoded into an image $x$. If $x$ is encoded into the latent space $z^{E}_{0}=E(x)$, its representation $z^{E}_{0}$ may not be aligned with the prompt used for inversion and generation (brown dashed line). As a result, the reconstructed $\hat{z}^{E}_{0}$ differs significantly from $z_{0}^{E}$, yielding an inconsistent image $x^{\prime}$. (b): Inconsistency can be fixed using a short noising and denoising iterations (red curves), which yields $\tilde{z}_{0}$. Then, applying the complete backward-forward process (green curves) can generate a consistent $\tilde{z}_{0}$.
How can this inconsistency be resolved? We propose to employ a short preprocessing step of fixed-point iteration [5]. The objective is to find a fixed point $z_{0}$ for the implicit function $z_{0}=h(z_{0},p)$, with $h$ representing the backward-forward process and $p$ is the prompt that would be used for inversion. The initial guess for the fixed-point iterations is $z^{E}_{0}=E(x)$. The iteration count should be kept small to prevent excessive alteration of the latent $z_{0}$ as this could potentially affect the visual appearance of the original image. In practical terms, we observed that a brief cycle of fixed-point iteration, involving two noise-addition steps (forward process) followed by two denoising steps with the prompt $p$ (backward process), proves effective.
This process is illustrated in Figure 4(b). Here, we pre-process $z^{E}_{0}$ by two fixed-point iterations to obtain an improved latent ${\tilde{z}}_{0}$ (red lines). ${\tilde{z}}_{0}$ is decoded to a very similar image $x$ and results with consistent inversion (dashed green lines). For 10k images, we found that the $L_{2}$ distance between $z_{0}$ and ${\tilde{z}}_{0}$ (i.e., $||z_{0}-{\tilde{z}}_{0}||$) is only $12$. This illustrates that the proposed pre-processing method is capable of maintaining a highly consistent backward-forward process. More detailed in Appendix F.
Editing with large guidance scale.
Table 1: Comparing inversion performance of iterative methods with and without prompt-aware adjustment, using large guidance scale.
PSNR #iter $T_{conv}$
[sec]
w/o Prompt-Aware Adjustment
Fixed-point [23] 10.2 diverge inf
Gradient Descent [13] 12.5 diverge inf
RNRI (ours) 19.3 3 2.8
w/ Prompt-Aware Adjustment
Fixed-Point 25.4 5 4.1
Gradient Descent 25.5 20 16.3
RNRI (ours) 25.8 2 2.61
Guidance scale plays a crucial role in editing tasks, with larger scales typically resulting in better and more natural edits [20]. Current iterative methods struggle to perform inversion with large guidance scales. This difficulty can be explained by the inconsistency between a given real image and a prompt, as described above. The guidance scale amplifies the influence of the prompt, intensifying this inconsistency and impeding convergence in current methods. [13] proposes the use of a different solver that requires significantly more iterations. [23] suggest inverting the image with a low guidance scale and then using a larger scale for editing specific regions identified by cross-attention maps. We demonstrate that prompt-aware adjustment enables iterative methods to perform inversion effectively by aligning a real image with a given prompt. Table 1 presents a comparison of inversion performance between different iterative methods, with and without prompt-aware adjustment. Guidance scale was set to $7.5$ (the default for StableDiffusion v2.1). The results indicate that prompt-aware adjustment enhances the performance of all methods in terms of PSNR, number of iterations (#iter), and convergence time ($T_{conv}$).
@@ 5 Experiments
We evaluate our approach on three main tasks: (1) Image inversion and reconstruction: Here, We assess the inversion fidelity by evaluating the quality of the reconstructed images. (2) Real-time Image Editing: We demonstrate the efficacy of our inversion scheme in image editing, highlighting its ability to facilitate real-time editing. (3) Rare Concept Generation: We illustrate how our method can be applied to improve the generation of rare concepts.
Compared Methods: We compared our approach with the following methods. (1) Standard DDIM Inversion [32]. (2) Null-text [22]. (3) EDICT [33]. (4) AIDI [23] (fixed-point based method). (5) ReNoise [10] (fixed-point based method) (6) ExactDPM [13] (gradient-based method). In all experiments, we used code and parameters provided by the respective authors.
Implementation details: To demonstrate the versatility of our approach, we conducted experiments on both the Latent Diffusion Model (LDM) using Stable Diffusion v2.1 and the Latent Consistency Model (LCM) using [30]. Input images were resized to 512 × 512. For LDM, we used 50 sampling steps, and for LCM, 4 steps. All baselines ran until convergence, while RNRI was run for 2 iterations per diffusion step. All methods were tested on a single A100 GPU for a fair comparison. PyTorch’s built-in gradient calculation was used for computing derivatives of Eq. 7.
Figure 5: (a) Inversion Results: Mean reconstruction quality (y-axis, PSNR) and runtime (x-axis, seconds) of four inversion methods on the COCO2017 validation set. (Left) Performance on latent diffusion model. All methods run for 50 inversion steps. Our method reaches high PSNR with ×$\times 4$ to ×$\times 12$ shorter inversion-reconstruction time compared to other methods. (Right) Performance on latent consistency model. All methods run for 2 inversion steps. Our method archives the highest PSNR in less than 0.5 seconds, significantly faster than other methods, which typically take 7 to 10 times longer. This allows for real-time inversion and editing capabilities using our approach. (b) Reconstruction qualitative results: Comparing image inversion-reconstruction performance. While DDIM fails to preserve a close connection to the original image, ReNoise creates rather a blurry image reflected in lower PSNR.
@@ 5.1 Image Reconstruction
To evaluate the fidelity of our approach, we measure PSNR of images reconstructed from seed inversions. Specifically, we used the entire set of 5000 images from the MS-COCO-2017 validation dataset [17], along with their corresponding captions. For each image-caption pair, we first found the inverted latent and then used it to reconstruct the image using the same caption. Given that the COCO dataset provides multiple captions for each image, we used the first listed caption as a conditioning prompt. Figure 5 (left) shows PSNR of reconstructed images in relation to inversion time, demonstrating the performance of our approach compared to SoTA inversion methods. Time is measured on a single NVIDIA A100 GPU for all methods. The dashed black line is the upper bound induced by the Stable Diffusion VAE. It shows that our method is able to achieve comparable PSNR to recent methods and close to the upper bound of VAE, yet accomplishes this in the shortest amount of time. Furthermore, in contrast to other methods, our approach accurately inverts an image to a single latent vector without requiring additional memory. Figure 5(right) further demonstrates a qualitative comparison of reconstructed images. We provide in the Appendix E an additional analysis.
Figure 6: Evaluation of editing performance: RNRI achieves superior CLIP and LPIPS scores, indicating better compliance with text prompts and higher structure preservation.
Figure 7: Qualitative results of image editing. RNRI edits images more naturally while preserving the structure of the original image. All baselines were executed until they reached convergence, whereas our approach was run for two iterations per diffusion step.
5.2 Real-Time Image Editing
@@
We now evaluate the effect of inverting images with RNRI, compared to state-of-the-art inversion baselines. We used Prompt2Prompt [11] as our editing method. We show that RNRI outperforms all baselines in editing performance while requiring the shortest time (less than 0.5 seconds per image, in latent consistency models). This opens the door for real-time editing capabilities. Videos where we show real-time editing performance using RNRI can be found in supplementary material (attached zip file). We present below both qualitative and quantitative results.
@@
Qualitative results. Figure 7 gives a qualitative comparison between SoTA approaches and RNRI in the context of real image editing. RNRI excels in accurately editing with target prompts producing images with both high fidelity to the input image and adherence to the target prompt. The examples illustrate how alternative approaches may struggle to retain the original structure or tend to overlook crucial components specified in the target prompt. As an example, in the second row of Figure 7, RNRI exclusively converts the kitten into a Lego figure, whereas other methods either fail to achieve this or alter the basket and branch as well. In the third row, all LDM-based methods struggle to accurately substitute bananas with oranges, and ReNoise [10] alters the background. In contrast, RNRI accurately edits the object while maintaining the original background. More results in supp. E.
@@
User Study. We further evaluated editing quality using human raters. We followed the evaluation process of [22] and asked human raters to rank edits made by three methods. 60 images were rated, 12 images were provided by the authors of [22] and the rest were randomly selected from the COCO dataset [17]. Three raters for each image were recruited through Amazon Mechanical Turk and were tasked to choose the image that better applies the requested text edit while preserving most of the original image. RNRI was preferred by raters in 40.4% of cases using Latent Diffusion Model (LDM), outperforming Null text  [22](12.2%), EDICT [34] (17.8%), and AIDI [23] (29.6%). For Latent Consistency Model (LCM), RNRI was overwhelmingly favored, with 89.9% preference compared to ReNoise’s [10] 10.1%.
LPIPS vs CLIP Score. Following [33, 15], we evaluate the results using two complementary metrics: LPIPS [35] to quantify the extent of structure preservation (lower is better) and a CLIP-based score to quantify how well the generated images comply with the text prompt (higher is better). Metrics are averaged across 100 MS-COCO images. Figure 6 illustrates that editing with RNRI yields a superior CLIP and LPIPS score, demonstrating the ability to perform state-of-the-art real image editing with superior fidelity. It further affirms the findings derived from the user study.
5.3 Seed Interpolation and Rare Concept Generation
Table 2: Image interpolation and centroid finding. In interpolation, two images $x^{1},x^{2}$ are inverted to generate images between their seeds $z_{T}^{1},z_{T}^{2}$. In centroid-finding, a set of images is inverted to find their centroid. Acc and FID scores improved using RNRI as the inversion method.
@@ 5.3 Seed Interpolation and Rare Concept Generation
AIDI [23] 52.01 6.13 68.14 5.32
RNRI (ours) 54.98 5.91 70.18 4.59
Table 3: Inversion Quality Impact on Rare Concept Generation: We assess image generation using a pre-trained classifier’s accuracy, comparing NAO [28] (with DDIM inversion), AIDI [23], and RNRI. We report average per-class accuracy for Head (over 1M samples), Medium, and Tail (rare classes <10K samples). RNRI enhances rare and medium concept accuracy without sacrificing overall performance.
ImageNet1k in LAION2B
Head Medium Tail Total Acc FID $\hat{T}_{Init}$ $\hat{T}_{Opt}$
n=235 n=509 n=256 (sec) (sec)
Methods #>1M 1M>#>10K 10K>#
DDIM inversion 98.5 96.9 85.1 94.3 6.4 25 29
AIDI 98.5 97.0 85.3 94.4 6.9 24 28
RNRI (ours) 98.6 97.9 89.1 95.8 6.3 17 25
Following the methodologies outlined in [29, 28], we attempt to generate images that are rare according to the diffusion training distribution. As demonstrated in these studies. This imbalance often leads to the generation of distorted or conceptually incorrect images. To address this issue, SeedSelect [29] takes a few images of a rare concept as input and uses a diffusion inversion module to iteratively refine the obtained seeds. These refined seeds are then used to generate new plausible images of the rare concept. NAO [28] extends this by introducing new paths and centroids for seed initialization. Both methods rely on DDIM Inversions, crucial for initial seed evaluation. Our work provides an alternative for precise inversion seeds, aiming for improved image quality and semantic accuracy. It’s important to highlight that most inversion techniques [22, 33, 15, 4] are not applicable in this context as they necessitate extra parameters for image reconstruction, which impedes the straightforward implementation of interpolation and centroid identification as suggested by NAO.
We now provide results for seed interpolation and rare concept generation.
@@
Interpolation and centroid finding: We evaluate RNRI for image interpolation and centroid finding. We follow the experimental protocol of [28] and compare images generated by DDIM with those produced by AIDI [23] and RNRI. Evaluation is conducted based on FID score and image accuracy, assessed using a pre-trained classifier. See details in [28]. Results are presented in Table 2. Notably, in comparison to DDIM and AIDI inversions, initializing with our seeds consistently results in higher-quality images both in interpolation paths and seed centroids.
@@
Rare concept generation: We further show the effect of our seed inversions on the performance of NAO centroids with SeedSelect for rare concept generation. Specifically, we compared images generated by SeedSelect initialized with NAO using DDIM inversion (NAO(DDIM) in Table 3) and using RNRI (NAO(RNRI)). We followed the evaluation protocol of [29, 28] on ImageNet1k classes arranged by their prevalence in the LAION2B dataset [31]. This dataset serves as a substantial “in the wild" dataset employed in training foundation diffusion models, such as Stable Diffusion [26]. Image quality is measured by FID and accuracy of a pre-trained classifier. For more details see [29, 28].
@@
The results, summarized in Table 3, demonstrate that our inversion method significantly boosts performance, both in accuracy and FID, compared to DDIM inversions. Furthermore, our inversions yield a more precise and effective initialization point for SeedSelect [29], resulting in notably quicker convergence (see Table 3) without compromising accuracy or image quality, in all categories (head, medium and tail) along with high gap at tail.
6 Limitations and Summary
Image inversion in diffusion models is vital for various applications like image editing, semantic augmentation, and generating rare concept images. Current methods often sacrifice inversion quality for computational efficiency, requiring significantly more compute resources for high-quality results. This paper presents Regularized Newton-Raphson Inversion (RNRI), a novel iterative approach that balances rapid convergence with superior accuracy, execution time, and memory efficiency. Using RNRI opens the door for real-time image editing. Despite its effectiveness, we encounter challenges. Despite demonstrating empirical convergence on a large-scale dataset, the model, like any iterative scheme, may fail to converge for certain images. We specifically identified such failure cases with incorrect prompts. (more details in Appendix G).
Supplemental Material
Appendix A Newton method for Multivariable Scalar Function
In this section, we will show the formulation of Newton’s method for zero crossing of a multi-variable scalar function. For a vector $\mathbf{x}\in\mathbb{R}^{n}$ and a function $f(\mathbf{x}):\mathbb{R}^{n}\rightarrow\mathbb{R}$ we are looking for the roots of the equation $f(\mathbf{x})=0$. Assuming that the function $f$ is differentiable, we can use Taylor expansion:
$f(\mathbf{x}+\delta)=f(\mathbf{x})+\nabla f~{}\delta^{T}+o(||\delta||^{2})$ (9)
For every $\delta,\mathbf{x}\in\mathbb{R}^{n}$. The idea of Newton’s method in higher dimensions is very similar to the one-dimensional scalar function; Given an iterate $\mathbf{x}^{k}$, we define the next iterate $\mathbf{x}^{k+1}$ by linearizing the equation $f(\mathbf{x}^{k})=0$ around $\mathbf{x}^{k}$ as above, and solving the linearized equation $f(\mathbf{x}^{k}+\delta)=0$. Dropping the higher orders in Equation 9 and solving the remaining linear equation for $\delta$ we get:
$\delta=-\frac{1}{n}\left[\frac{1}{\frac{\partial f}{\partial x_{1}}},\frac{1}{\frac{\partial f}{\partial x_{2}}},...,\frac{1}{\frac{\partial f}{\partial x_{n}}}\right]f(\mathbf{x}^{k})$ (10)
where $x_{i}$ indicates the $i$-th component of vector $\mathbf{x}$. The above can be easily proven by substitution of $\delta$ into Equation 9. Using the relation, $\delta=\mathbf{x}^{k+1}-\mathbf{x}^{k}$ we get the final iterative scheme:
$x_{i}^{k+1}=x_{i}^{k}-\frac{1}{n}\frac{f(\mathbf{x}^{k})}{\frac{\partial f}{\partial x_{i}}(\mathbf{x}^{k})}$ (11)
Note that this equation is component-wise i.e. each component is updated separately, facilitating the solution.
Appendix B Seed Exploration & Rare concept generation